OpenStudy (anonymous):

I will medal and fan if you help, The range of the function f(k) = k2 + 2k + 1 is {25, 64}. What is the function’s domain?

OpenStudy (anonymous):

@ajo21

OpenStudy (anonymous):

{5, 8} {-5, -8} {3, 8} {4, 7} {4, 8} these are the answers

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (pawanyadav):

K2+2k+1=(k+1)^2 So when range (25,64) Then.. K+1=5 or k+1=8 K=4 or k=7 Do domain (4,7)

OpenStudy (anonymous):

thanks

OpenStudy (pawanyadav):

Welcome